# Assessing uncertainty in Gaussian mixtures-based entropy estimation

### v1

@@ Abstract
Entropy estimation plays a crucial role in various fields, such as information theory, statistical data science, and machine learning. However, traditional entropy estimation methods often struggle with complex data distributions. Mixture-based estimation of entropy has been recently proposed and gained attention due to its ease of use and accuracy. This paper presents a novel approach to quantify the uncertainty associated with this mixture-based entropy estimation method using weighted likelihood bootstrap. Unlike standard methods, our approach leverages the underlying mixture structure by assigning random weights to observations in a weighted likelihood bootstrap procedure, leading to more accurate uncertainty estimation. The generation of weights is also investigated, leading to the proposal of using weights obtained from a Dirichlet distribution with parameter $\alpha=0.8137$ instead of the usual $\alpha=1$. Furthermore, the use of centered percentile intervals emerges as the preferred choice to ensure empirical coverage close to the nominal level. Extensive simulation studies comparing different resampling strategies are presented and results discussed. The proposed approach is illustrated by analyzing the log-returns of daily Gold prices at COMEX for the years 2014–2022, and the Net Rating scores, an advanced statistic used in basketball analytics, for NBA teams with reference to the 2022/23 regular season.
@@ 1 Introduction
Closed-form expressions for the entropy of several univariate distributions are available in the literature (Michalowicz et al.,, 2014), but only few multivariate distributions admit a solution in analytic form. For parametric pdfs $\{f(\boldsymbol{x};\boldsymbol{\theta}):\boldsymbol{\theta}\in\boldsymbol{\Theta}\}$, when a closed-form expression is available, using the MLE $\widehat{\boldsymbol{\theta}}$ guarantees that the estimate of the entropy is both asymptotically unbiased and efficient (Kay,, 1993, Theorem 7.3). If we cannot assume a specific parametric distribution, an estimate of the entropy is often obtained using nonparametric estimation of $f(\boldsymbol{x})$ based on histogram, kernel density, or nearest neighbour estimators. This latter approach is, however, limited to low-dimensional cases, often $d\leq 2$. Recently, a simple and intuitive mixture-based estimator has been proposed by Robin and Scrucca, (2023) using finite mixture models (FMMs). The authors showed with extensive simulation studies the accuracy and efficiency of the proposal, which compares favorably against other estimators often used in practice.
The remainder of this paper is structured as follows: Section 2 reviews the mixture-based approach for entropy estimation and its advantages over traditional methods. Section 3 introduces our proposal for quantifying uncertainty in mixture-based entropy estimation. Specifically, we propose to adopt a weighted likelihood bootstrap (WLB) approach by assigning random weights drawn from a Dirichlet distribution when refitting mixture models. Generation of weights is discussed in Section 4, where we introduce a novel proposal instead of the usually adopted uniform Dirichlet over the simplex. Extensive simulation studies are presented in Section 5, where the performance of different resampling strategies and percentile bootstrap intervals are compared. Section 6 illustrate the methodology by analyzing entropy in two real-world datasets. The final section concludes with a summary of findings and potential future research directions.
@@ 2 Mixture-based entropy estimation
where $\boldsymbol{\theta}=\{\boldsymbol{\pi},\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{G}\}$ are the unknown parameters, with $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{G})$ the mixing weights (s.t. $\pi_{k}>0,\sum_{k=1}^{G}\pi_{k}=1)$, and $\boldsymbol{\theta}_{k}$ the parameters associated to the density $f_{k}(\cdot)$ of the $k$th component of the mixture ($k=1,\dots,G$). Often, the densities $f_{k}(\cdot)$ are taken to belong to the same parametric family of distributions, but with different parameters $\boldsymbol{\theta}_{k}$.
@@
Recalling that, from a generative point of view, FMMs in (1) can be re-expressed as a hierarchical model by introducing the multinomial discrete latent variable $Z\in\{1,\dots,G\}$ and writing
@@
$X\mid(Z=k)$ $\sim f_{k}(\boldsymbol{x};\boldsymbol{\theta}_{k}),$
then by the law of total expectation, the entropy can be written as
@@
Based on this formulation, (Robin and Scrucca,, 2023) showed that a simple closed-formula mixture-based estimator for samples of size $n$ can be calculated as follows:
@@
where $\boldsymbol{\theta}=\{\boldsymbol{\pi},\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{G},\boldsymbol{\Sigma}_{1},\dots,\boldsymbol{\Sigma}_{G}\}$ are the unknown parameters of the mixture, with $\boldsymbol{\pi}=(\pi_{1},\pi_{2},\dots,\pi_{G})$ the mixing weights, and $\phi(\boldsymbol{x};\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ the underlying multivariate Gaussian pdf of the $k$th component with mean vector $\boldsymbol{\mu}_{k}$ and covariance matrix $\boldsymbol{\Sigma}_{k}$. The entropy estimate from (2) can then be written as
@@
where $(\widehat{\pi}_{k},\widehat{\boldsymbol{\mu}}_{k},\widehat{\boldsymbol{\Sigma}}_{k})$ are the estimated parameters of $k$th component of the GMM obtained using the EM algorithm. For details see Scrucca et al., (2023, Sec. 2.2).
@@ 3 Uncertainty assessment via weighted likelihood bootstrap
Uncertainty assessment in entropy estimation is crucial for reliable interpretation of estimates. However, existing literature lacks methodologies specifically designed for this purpose. To fill this gap, we propose adopting a resampling approach based on the bootstrap (Efron and Tibshirani,, 1993; Davison and Hinkley,, 1997) to quantify uncertainty effectively.
Standard nonparametric bootstrap (BS) is a resampling technique that relies on repeatedly drawing samples with replacement from the observed data (Efron,, 1982). This procedure is equivalent to the repeated generation of random weights from a multinomial distribution with equal probabilities, i.e. $\mathcal{M}(1;1/n,...,1/n)$. In contrast, Weighted Likelihood Bootstrap (WLB) is a resampling procedure that generate bootstrap samples by assigning random weights to observations generated from a Dirichlet distribution over the $(n-1)$-simplex, i.e.
@@
$\ell_{C}^{W}(\boldsymbol{\theta})=\sum_{i=1}^{n}\sum_{k=1}^{G}z_{ig}\,w_{i}\log\left\{\pi_{k}\,\phi(\boldsymbol{x};\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})\right\},$ (5)
where $\boldsymbol{\theta}=(\pi_{1},\dots,\pi_{G-1},\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{G},\boldsymbol{\Sigma}_{1},\dots,\boldsymbol{\Sigma}_{G})$ are the parameters to be estimated.
Let $(\mathcal{M},G)$ be the Gaussian mixture model $\mathcal{M}$ defined by a parsimonious covariance decomposition (Scrucca et al.,, 2023, Sec. 2.2.1) with $G$ number of mixture components. These represent tuning parameters that can be set by the researcher or determined using a model selection criterion, such as the Bayesian information criterion (Schwarz,, 1978, BIC) or the integrated complete-data likelihood criterion (Biernacki et al.,, 2000, ICL).
The proposed WLB procedure for a $(\mathcal{M},G)$ can be described as follows:
@@
    fit $(\mathcal{M},G)$ using the data $\{\boldsymbol{x}_{i},w_{i}\}_{i=1}^{n}$ by maximizing (5) to get the bootstrap estimate of parameters $\widehat{\boldsymbol{\theta}}^{\,*}_{b}=\left\{\widehat{\pi}^{\,*}_{k},\widehat{\boldsymbol{\mu}}^{\,*}_{k},\widehat{\boldsymbol{\Sigma}}^{\,*}_{k}\right\}_{k=1}^{G}$;
@@
    replicate steps 1–3 a large number of times, say $B$, to obtain the WLB distribution $\left\{\widehat{H}^{*}_{b}(X)\right\}_{b=1}^{B}$ as an approximate posterior sampling.
@@
$\left[\widehat{H}^{*}_{\alpha/2},\;\widehat{H}^{*}_{1-\alpha/2}\right],$ (6)
@@
$\Pr\left(\widehat{H}^{*}_{\alpha/2}-\widehat{H}(X)\leq\widehat{H}(X)-H(X)\leq\widehat{H}^{*}_{1-\alpha/2}-\widehat{H}(X)\right)\approx 1-\alpha$
$\Pr\left(2\widehat{H}(X)-\widehat{H}^{*}_{1-\alpha/2}\leq H(X)\leq 2\widehat{H}(X)+\widehat{H}^{*}_{\alpha/2}\right)\approx 1-\alpha.$
Notice that, if $\widehat{H}^{*}_{b}(X)-\widehat{H}(X)$ is the bias for the $b$th bootstrap sample, then a bias-corrected estimate is given by
@@
Thus, a WLB centered percentile interval at the level $(1-\alpha)100\%$ can be computed using the percentiles from the distribution of bias-corrected entropy estimates $\left\{\widehat{H}^{*c}_{b}(X)\right\}_{b=1}^{B}$, i.e. by computing
@@ 4 On the generation of weights in WLB
In WLB, as well as in BB and WBB methods, the weights associated with each observation are generated from a Dirichlet distribution, i.e. $\boldsymbol{w}\sim\mathcal{D}(\boldsymbol{\alpha})$ with $\boldsymbol{\alpha}=\alpha\boldsymbol{1}_{n}$, where $\boldsymbol{1}_{n}$ is a vector of ones of length $n$. In general, generating random values from a $\mathcal{D}(\boldsymbol{\alpha})$ distribution is equivalent to generating them from independent Gamma distributions $\mathcal{G}(\alpha,\beta=1)$, resulting in both the expected value and the variance being equal to $\alpha$. It’s worth noting that when $\alpha=1$, an Exponential distribution with a rate parameter of 1 is obtained.
@@
Here, we propose using $\alpha=0.8137$, a value derived by mimicking a well-known result in standard nonparametric bootstrap. Recalling that the probability of an observation being included at least once in a nonparametric bootstrap sample is $1-\exp(-1)\approx 0.6321$, we aim to determine the value of $\alpha$ that ensures, on average, a median weight of $0.6321$ is assigned to each observation.
@@ 5.2 Data generation settings
Data were simulated from the aforementioned distributions for sample sizes $n=\{100,200,500,1000\}$. For each setting, 1,000 replications were generated. A GMM was then fitted to each synthetic dataset using the mclust R package (Fraley et al.,, 2024), with number of mixture components and parsimonious variances (or covariance matrices in case of multidimensional data) selected using the Bayesian information criterion (BIC Schwarz,, 1978).
@@ 5.3 Results
Tables 1 and 37 in A present, for each investigated distribution, summaries of the simulation results for the bootstrap methods under comparison. For the same settings, the associated Figures 1 and 1014 in A show the empirical coverage of different bootstrap intervals (panel a) and the ratio of bootstrap absolute bias over the bootstrap standard error (panel b).
@@
Table 1: Simulation results for data generated from a $t$ distribution with $\text{df}=3$ degrees of freedom. For each bootstrap method and sample size, the average estimate of entropy, the bootstrap estimates of bias and standard error, and the associated bootstrap percentile intervals are reported.
@@
Figure 2: Simulation results for data generated from a $t$ distribution with $\text{df}=3$ degrees of freedom. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weigthed likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.

### v2

@@ Abstract
Entropy estimation plays a crucial role in various fields, such as information theory, statistical data science, and machine learning. However, traditional entropy estimation methods often struggle with complex data distributions. Mixture-based estimation of entropy has been recently proposed and gained attention due to its ease of use and accuracy. This paper presents a novel approach to quantify the uncertainty associated with this mixture-based entropy estimation method using weighted likelihood bootstrap. Unlike standard methods, our approach leverages the underlying mixture structure by assigning random weights to observations in a weighted likelihood bootstrap procedure, leading to more accurate uncertainty estimation. The generation of weights is also investigated, leading to the proposal of using weights obtained from a specific Dirichlet distribution which, in conjunction with centered percentile intervals, yields the optimal setting to ensure empirical coverage closer to the nominal level. Extensive simulation studies comparing different resampling strategies are presented and results discussed. The proposed approach is illustrated by analyzing the log-returns of daily Gold prices at COMEX for the years 2014–2022, and the Net Rating scores, an advanced statistic used in basketball analytics, for NBA teams with reference to the 2022/23 regular season.
@@ 1 Introduction
Closed-form expressions for the entropy of several univariate distributions are available in the literature (Michalowicz et al.,, 2014), but only few multivariate distributions admit a solution in analytic form. For parametric pdfs $\{f(\boldsymbol{x};\boldsymbol{\theta}):\boldsymbol{\theta}\in\boldsymbol{\Theta}\}$, when a closed-form expression is available, using the maximum likelihood estimate (MLE) $\widehat{\boldsymbol{\theta}}$ guarantees that the estimate of the entropy is both asymptotically unbiased and efficient (Kay,, 1993, Theorem 7.3). If we cannot assume a specific parametric distribution, an estimate of the entropy is often obtained using nonparametric estimation of $f(\boldsymbol{x})$ based on histogram, kernel density, or nearest neighbour estimators. This latter approach is, however, limited to low-dimensional cases, often $d\leq 2$. Recently, a simple and intuitive mixture-based estimator has been proposed by Robin and Scrucca, (2023) using finite mixture models (FMMs). The authors showed with extensive simulation studies the accuracy and efficiency of the proposal, which compares favorably against other estimators often used in practice.
The remainder of this paper is structured as follows: Section 2 reviews the mixture-based approach for entropy estimation and its advantages over traditional methods. Section 3 introduces our proposal for quantifying uncertainty in mixture-based entropy estimation. Specifically, we propose to adopt a weighted likelihood bootstrap (WLB) approach by assigning random weights drawn from a Dirichlet distribution when refitting mixture models. Generation of weights is discussed in Section 4, where we introduce a novel proposal instead of the usually adopted uniform Dirichlet over the simplex. Extensive simulation studies are presented in Section 5, where the performance of different resampling strategies and percentile bootstrap intervals are compared. Section 6 illustrates the methodology by analyzing entropy in two real-world datasets. The final section concludes with a summary of findings and potential future research directions.
@@ 2 Mixture-based entropy estimation
where $\boldsymbol{\theta}=\{\boldsymbol{\pi},\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{G}\}$ are the unknown parameters, with $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{G})$ the mixing weights (s.t. $\pi_{k}>0,\sum_{k=1}^{G}\pi_{k}=1)$, and $\boldsymbol{\theta}_{k}$ the parameters associated to the density $f_{k}(\cdot)$ of the $k$-th component of the mixture ($k=1,\dots,G$). Often, the densities $f_{k}(\cdot)$ are taken to belong to the same parametric family of distributions, but with different parameters $\boldsymbol{\theta}_{k}$.
@@
Recalling that, from a generative point of view, FMMs in (1) can be re-expressed as a hierarchical model by introducing the multinomial discrete latent variable $Z=(Z_{1},\dots,Z_{G}){}^{\!\top}$, where $Z_{k}=1$ if the component of origin of $X$ in the mixture is equal to $k$ and 0 otherwise ($k=1,\dots,G$), and assuming
@@
$X\mid(Z_{k}=1)$ $\sim f_{k}(\boldsymbol{x};\boldsymbol{\theta}_{k}),$
where $\mathcal{M}(1,\boldsymbol{\pi})$ is the multinomial distribution consisting of one draw on $G$ categories with probabilities $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{G}){}^{\!\top}$. Then, by the law of total expectation, the entropy can be written as
@@
Based on this formulation, Robin and Scrucca, (2023) showed that a simple closed-formula mixture-based estimator for samples of size $n$ can be calculated as follows:
@@
where $\boldsymbol{\theta}=\{\boldsymbol{\pi},\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{G},\boldsymbol{\Sigma}_{1},\dots,\boldsymbol{\Sigma}_{G}\}$ are the unknown parameters of the mixture, with $\boldsymbol{\pi}=(\pi_{1},\pi_{2},\dots,\pi_{G})$ the mixing weights, and $\phi(\boldsymbol{x};\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ the underlying multivariate Gaussian pdf of the $k$-th component with mean vector $\boldsymbol{\mu}_{k}$ and covariance matrix $\boldsymbol{\Sigma}_{k}$. The entropy estimate in (2) can then be written as
@@
where $(\widehat{\pi}_{k},\widehat{\boldsymbol{\mu}}_{k},\widehat{\boldsymbol{\Sigma}}_{k})$ are the estimated parameters of $k$-th component of the GMM obtained using the EM algorithm. For details see Scrucca et al., (2023, Sec. 2.2).
@@ 3 Uncertainty assessment via weighted likelihood bootstrap
Uncertainty assessment in entropy estimation is crucial for reliable interpretation of estimates. Although some authors have addressed this issue in the context of entropy for discrete distributions (Amiri,, 2014; Pesenti et al.,, 2017), existing literature lacks methodologies specifically designed for differential entropy. To fill this gap, we propose adopting a resampling approach based on the bootstrap (Efron and Tibshirani,, 1993; Davison and Hinkley,, 1997) to quantify uncertainty effectively.
Standard nonparametric bootstrap (BS) is a resampling technique that relies on repeatedly drawing samples with replacement from the observed data (Efron,, 1982). This procedure is equivalent to the repeated generation of random weights from a multinomial distribution with equal probabilities, i.e. $\mathcal{M}(1;\;\tfrac{1}{n}\boldsymbol{1}_{n})$, where $\boldsymbol{1}_{n}$ is a vector of ones of length $n$. In contrast, Weighted Likelihood Bootstrap (WLB) is a resampling procedure that generate bootstrap samples by assigning random weights to observations generated from a Dirichlet distribution over the $(n-1)$-simplex, i.e.
@@
$\ell_{C}^{W}(\boldsymbol{\theta})=\sum_{i=1}^{n}\sum_{k=1}^{G}z_{ik}\,w_{i}\log\left\{\pi_{k}\,\phi(\boldsymbol{x}_{i};\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})\right\},$ (5)
where $\boldsymbol{\theta}=(\pi_{1},\dots,\pi_{G-1},\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{G},\boldsymbol{\Sigma}_{1},\dots,\boldsymbol{\Sigma}_{G})$ are the parameters to be estimated, and $z_{ik}=\pi_{k}\phi(\boldsymbol{x}_{i};\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})/\sum_{g=1}^{G}\pi_{g}\phi(\boldsymbol{x}_{i};\boldsymbol{\mu}_{g},\boldsymbol{\Sigma}_{g})$, the conditional probability that the $i$-th observation arose from mixture component $k$.
Let $(M,G)$ be the Gaussian mixture model $M$ defined by a parsimonious covariance decomposition (Scrucca et al.,, 2023, Sec. 2.2.1) with $G$ number of mixture components. These represent tuning parameters that can be set by the researcher or determined using a model selection criterion, such as the Bayesian information criterion (BIC; Schwarz,, 1978) or the integrated complete-data likelihood criterion (ICL; Biernacki et al.,, 2000).
The proposed WLB procedure for a $(M,G)$ can be described as follows:
@@
    fit $(M,G)$ using the data $\{\boldsymbol{x}_{i},w_{i}\}_{i=1}^{n}$ by maximizing (5) to get the bootstrap estimate of parameters $\widehat{\boldsymbol{\theta}}^{\,*}_{b}=\left\{\widehat{\pi}^{\,*}_{k},\widehat{\boldsymbol{\mu}}^{\,*}_{k},\widehat{\boldsymbol{\Sigma}}^{\,*}_{k}\right\}_{k=1}^{G}$;
@@
    replicate steps 1–3 a large number of times, say $B$, to obtain the WLB distribution $\widehat{H}^{*}(X)=\left\{\widehat{H}^{*}_{b}(X)\right\}_{b=1}^{B}$ as an approximate posterior sampling.
@@
$\left[\widehat{H}^{*}(X;\alpha/2),\;\widehat{H}^{*}(X;1-\alpha/2)\right],$ (6)
@@
$\Pr\left(\widehat{H}^{*}(X;\alpha/2)-\widehat{H}(X)\leq\widehat{H}(X)-H(X)\leq\widehat{H}^{*}(X;1-\alpha/2)-\widehat{H}(X)\right)\approx 1-\alpha$
$\Pr\left(2\widehat{H}(X)-\widehat{H}^{*}(X;1-\alpha/2)\leq H(X)\leq 2\widehat{H}(X)-\widehat{H}^{*}(X;\alpha/2)\right)\approx 1-\alpha.$
Notice that, if $\widehat{H}^{*}_{b}(X)-\widehat{H}(X)$ is the bias for the $b$-th bootstrap sample, then a bias-corrected estimate is given by
@@
Thus, a WLB centered percentile interval at the level $(1-\alpha)100\%$ can be computed using the percentiles from the distribution of bias-corrected entropy estimates $\widehat{H}^{*c}(X)=\left\{\widehat{H}^{*c}_{b}(X)\right\}_{b=1}^{B}$, i.e. by computing
@@ 4 On the generation of weights in WLB
In WLB, as well as in BB and WBB methods, the weights associated with each observation are generated from a Dirichlet distribution, i.e. $\boldsymbol{w}\sim\mathcal{D}(\boldsymbol{\alpha})$ with $\boldsymbol{\alpha}=\alpha\boldsymbol{1}_{n}$. In general, generating random values from a $\mathcal{D}(\boldsymbol{\alpha})$ distribution is equivalent to generating them from independent Gamma distributions $\mathcal{G}(\alpha,\beta=1)$, resulting in both the expected value and the variance being equal to $\alpha$. It is worth noting that when $\alpha=1$, an Exponential distribution with a rate parameter of 1 is obtained.
@@
Here, we propose using $\alpha=0.8137$, a value derived by mimicking a well-known result in standard nonparametric bootstrap. Recalling that the probability of an observation being included at least once in a nonparametric bootstrap sample is $1-\exp(-1)\approx 0.6321$ as $n\to\infty$ (Efron and Tibshirani,, 1993, Problem 17.7), we aim to determine the value of $\alpha$ that ensures, on average, a median weight of $0.6321$ is assigned to each observation.
@@ 5.2 Data generation settings
Data were simulated from the aforementioned distributions for sample sizes $n=\{100,200,500,1000\}$. For each setting, 1,000 replications were generated. A GMM was then fitted to each synthetic dataset using the mclust R package (Fraley et al.,, 2024), with number of mixture components and parsimonious variances (or covariance matrices in case of multidimensional data) selected using the Bayesian information criterion (BIC; Schwarz,, 1978).
@@ 5.3 Results
Tables 1 and 37 in Appendix A present, for each investigated distribution, summaries of the simulation results for the bootstrap methods under comparison. For the same settings, the associated Figures 2 and 913 in Appendix A show the empirical coverage of different bootstrap intervals (panel a) and the ratio of bootstrap absolute bias over the bootstrap standard error (panel b).
@@
Table 1: Simulation results for data generated from a $t$ distribution with $\text{df}=3$ degrees of freedom. For each bootstrap method and sample size, the table reports averaged values of the entropy estimates, bootstrap estimates of bias, standard error, and percentile interval limits, along with their empirical coverage.
@@
Figure 2: Simulation results for data generated from a $t$ distribution with $\text{df}=3$ degrees of freedom. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weighted likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.